OpenStudy (anonymous):

I need help with solving roots. I can't understand it, please help. Problems written in comments, medals for best answer!

OpenStudy (anonymous):

\[\sqrt{18}\]

OpenStudy (anonymous):

I know that the square root of 16 is 4, but I can't understand these.

OpenStudy (loser66):

18=2*9 =2*3^2

OpenStudy (dan815):

umm sqrt of 16 is -4 and 4

OpenStudy (dan815):

because -4 * -4 = 16 and 4* 4 = 16

OpenStudy (anonymous):

I know the root of 16, but I don't know how to solve 18.

OpenStudy (anonymous):

@Loser66 can you please explain your example?

OpenStudy (loser66):



OpenStudy (dan815):